OpenStudy (bossimbacon):

Which expressions are equivalent to 4^3 x 4^-5? Choose all answers that are correct. A) 4^-2 B) 1/16 C) 4^-15 D) 0.0625 E) 16^-1 F) 1/4^15 G) 16^-2 H) 1/4^2 I) (4^-1)^2 J) 4^3/4^5

OpenStudy (bossimbacon):

@robtobey

OpenStudy (anonymous):

\[\frac{4^3}{4^5}=\frac{1}{4^2} \]

OpenStudy (bossimbacon):

so whats the answer o.O

OpenStudy (bossimbacon):

i don't get it :/